A triangle has an angle that measures 60°. The other two angles are in a ratio of 5:7. What are the measures of those two angles?

Answer by ikleyn(52800)

.

The two other angles are 5x and 7x, and

    5x + 7x = 180 - 60 degs.


Simplify and find x

     12x   = 120

       x   = 120/12 = 10.


ANSWER.  Two other angles are 50 degs and 70 degs.

Solved.